# Integrating large language models and active inference to understand eye movements in reading and dyslexia

### v2

@@ 2.2 Simulation 2: Reading Sentences
This task involves reading and recognizing 100 sentences composed of 4 or 8 words, with each word consisting of 1 to 4 syllables, generated by BERT (Devlin et al., 2019). The sentences with 4 words have an average length of $25.96\pm 1.96$ letters within the range of $[21,31]$. Similarly, sentences with 8 words have an average length of $50.24\pm 1.64$ letters, in the interval of $[47,55]$ (see Table S.13). To add complexity, we carefully selected sentences with a substantial word overlap (Table S.15). For this task, we utilize the comprehensive model depicted in Figure 1, encompassing three hierarchical levels.
Previous studies have shown that proficient readers can scan lines of text using only a few saccades, while dyslexic individuals exhibit an increased number of shorter saccades (Hutzler and Wimmer, 2004; Rayner, 1983; De Luca et al., 1999) and fewer word skipping occurrences (Bucci et al., 2008; Jainta and Kapoula, 2011; Hawelka et al., 2010) compared to controls. Additionally, their performance decrease and increased number of saccades are influenced by the number of words in the text (see (De Luca et al., 1999), Fig. 2). To verify whether our model accurately replicates these findings, we compare a Control model (CM) with no noise and a Dyslexic model (DM) with noise introduced in the transition functions of all three hierarchical levels.
Figure 7: Simulation 2 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at both hierarchical levels, and three alternative versions of the Dyslexic Model, where we added noise only at the level of syllables (DM - Noise on Level 1), words (DM - Noise on Level 2), or sentences (DM - Noise on Level 3).
@@
Collectively, these results build upon those observed for single words (Simulation 1) and successfully replicate empirical findings on dyslexic individuals’ reading of sentences. The trends illustrated in Figure 7(a) closely resemble the outcomes reported (De Luca et al., 1999, Fig. 3(A)), which demonstrates that dyslexics make more saccades compared to controls while reading single lines of text (equivalent to reading 8-word sentences in Figure 7(a)). Moreover, the number of saccades exhibited by the DM in our simulations aligns remarkably well with the numerical outcomes from the same study. Similarly, the outcomes in Figure 8(a) closely resemble the findings depicted in (De Luca et al., 1999, Fig. 3(B)), revealing that dyslexics make shorter forward saccades than controls. In this context, the amplitudes of both dyslexics and the DM model closely correspond.
Figure 8: Results of Simulation 2: amplitude of forward (a) and backward (b) saccades when reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise and the Dyslexic Model (DM) with noise at all the hierarchical levels.
In summary, our results effectively reproduce the observation that dyslexic individuals make a greater number of shorter saccades while reading sentences, and their reading performance is dependent on the number of words in the sentence (refer to (De Luca et al., 1999, Fig.2)). The causes of these impairments are akin to those discussed for single words (Simulation 1), but their effects become more pronounced as sentence comprehension necessitates the integration of information over more extended periods. For instance, the high occurrence of ample backward saccades is a result of the DM frequently losing context and subsequently needing to backtrack to the initial words of a sentence to regain it. This further clarifies why noise at the sentence level substantially impairs this task (Figure 7).
@@
Up to this point, our simulations involved reading known words and sentences, already present in the BERT vocabulary. Hence, the model could confidently assign probabilities at the second and third layers. Nevertheless, our model can also read unknown words and sentences, despite its inability to assign them a prior probability. When confronted with unknown words or sentences, the model reads syllable-by-syllable (or word-by-word) and simultaneously acknowledges uncertainty at the second (or third) level. This mode aligns with the nonlexical (or sublexical) route in dual-route theories of reading (Coltheart, 2005).
@@
To demonstrate this capability, we simulate reading the last four words of the sentence This paper is also framed in an offbeat manner and compare two versions of the model: one in which all the words exist in the model’s vocabulary (Figure 9(a)) and another where the word offbeat is absent from the vocabulary (Figure 9(b)). In the version with all known words, the model behaves similarly to the previous simulations, accurately recognizing all the words and the sentence. This is evident in Figure 9(a), where probabilities at the second and third levels quickly converge to one after a few iterations.
@@
Conversely, the version without knowledge of the word offbeat repeatedly makes saccades to the syllables of this word (at Level 1) until it reaches the maximum number of iterations set for this simulation. It correctly recognizes each syllable and can read the word aloud, potentially adding it to its vocabulary (Coltheart, 2005). However, it does not recognize the word (at Level 2) or the sentence (at Level 3). This is evident in Figure 9(b), where probabilities at the second and third levels never exceed 0.5.
@@
The example in Fig. 9 illustrates the distinction between reading a known versus a novel word and the termination conditions for inference at each hierarchical level. The inference can terminate in two ways: first, when the model reaches a threshold of confidence (entropy of hidden states, $\chi^{(i)}$) about the to-be-recognized syllable, word, or sentence. This happens when the model successfully infers a known syllable, word, or sentence.
@@
As shown in Figure 9, Level 1 successfully infers all the syllables within 3 iterations and never reaches $K_{\max}^{(1)}$. However, Level 2 fails to recognize the unknown word offbeat: it continues jumping between the syllables off and beat until it reaches the maximum number of allowed iterations (6 iterations, indicated in the grey panel). When the Level 2 inference of the current word is halted, a low confidence message is reported to the level above and the inference continues with the next words (that are successfully recognized). Similarly, Level 3 fails to infer the unknown sentence after the maximum number of iterations (7 iterations), despite correctly recognizing most of the words. At the end of the simulation, the model identifies that the cause of failure is at Level 2, as Level 1 correctly recognized all the syllables. Consequently, the model reads the unknown word syllable-by-syllable, i.e. using the nonlexical route in dual route theories of reading (Coltheart, 2005).
Figure 9: Simulation 3 results: Reading known vs. unknown words. Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time, while reading (a) known words or (b) the unknown word offbeat. The blue circles at the bottom level indicate saccades. The figure format is consistent with Figure 1(d).
@@ 2.4 Simulation 4: Knowing the topic improves sentence recognition
To illustrate this, we conduct a task involving the recognition of 100 sentences, each containing 9 words, generated by BERT. These sentences comprise an average of $69.55\pm 5.02$ letters, ranging between 61 and 82 characters (see Table S.13). We allocate one-third of the sentences to each of the three topics of the European Research Council (ERC): Physical Sciences and Engineering (PE), Life Sciences (LS), and Social Sciences and Humanities (SH).
We compare the performance of two versions of the model: one with a flat prior (33.3%) regarding the sentence being read, and another with an informative prior that assigns a probability of 100% to the correct topic. As depicted in Figure  10, the model with an informative prior requires significantly fewer total saccades and backward saccades than the model with a flat prior (see Table S.11 and Table S.12 for statistical comparisons). This advantage arises because knowing the topic allows the model to predict certain words with high confidence, enabling it to skip unnecessary saccades during the reading process. Intriguingly, despite the model with an informative prior skipping several words, it does not necessitate more backward (corrective) saccades compared to the model with a flat prior. This indicates that the prior information promotes a better speed-accuracy trade-off.
Figure 10: Simulation 4 results: How prior information affects the reading process. The figure displays the total saccades (a) and backward saccades (b) of two model versions: one with a flat prior about the topic of the text (flat prior) and another with certain knowledge of the topic (informative prior). The comparison highlights the impact of prior information on the number and nature of saccades during reading.
@@ 3 Discussion
In this paper, we presented a novel hierarchical active inference model of reading, which combines the robust generative abilities of large language models (Devlin et al., 2019; Radford et al., 2019) with the inferential and information-gathering capacities of active inference (Friston, 2010; Parr et al., 2022).
@@
Large language models, like BERT (Devlin et al., 2019) and GPT (Generative Pre-trained Transformer) (Radford et al., 2019), have demonstrated remarkable performance in various natural language processing tasks, including question answering, language translation, text classification, and sentence prediction. These models can learn essential linguistic structures without external supervision, making them valuable tools to explore language processing in the brain (Manning et al., 2020; Goldstein et al., 2022). However, they lack the active strategies employed by human readers, such as the ability to make saccades selectively on the most informative parts of the text (Ferro et al., 2010; Norris, 2006).
@@
To capture active reading, we integrate BERT (Devlin et al., 2019) into a hierarchical active inference framework (Friston, 2010; Parr et al., 2022). Active inference is a theoretical framework that describes how the brain employs probabilistic inference over a generative model to effectively sample sensory inputs, minimizing variational free energy or prediction errors. For modeling reading, we utilize a hierarchical generative model, aligning with evidence suggesting that language processing in the brain operates through hierarchically organized predictions and prediction errors (Schmitt et al., 2021; Caucheteux et al., 2023; Heilbron et al., 2022).
@@
By placing BERT at the top of the hierarchy, we ensure accurate next-word predictions and the model’s capability to tackle real-world reading tasks. Beyond next-word prediction, our hierarchical model facilitates both inferring the content being read and predicting future elements across different linguistic levels, such as syllables, words, and sentences. Our simulation results demonstrate the model’s perfect accuracy in inferring words (Simulation 1) and sentences (Simulation 2) during reading. Additionally, this hierarchical structure enables the model to read novel words not present in BERT’s vocabulary (Simulation 3), assembling them syllable-by-syllable, representing the nonlexical (or sublexical) route in dual route theories of reading (Coltheart, 2005). Moreover, the model can incorporate prior knowledge, like the topic of the sentence, to expedite the reading process (Simulation 4). Significantly, the model offers interpretability through its alignment of distinct levels with established linguistic constituents (sentences, words, syllables), which could correspond to distributed neural populations or cell assemblies within the brain (Heilbron et al., 2022; Pulvermüller, 1999). While delving into the neural mechanisms that underlie this proposed framework extends beyond the confines of this article, future studies could encompass the simulation of neuronal dynamics during the inference process, in terms of variational free energy minimization (Friston et al., 2017; Isomura et al., 2022).
@@
Using active inference, our model characterizes reading as an active, hypothesis testing process, allowing us to simulate eye movements during reading. The core idea is that saccades are directed to the most informative parts of the text, enabling the model to test its predictions and reduce uncertainty about the text being read (Ferro et al., 2010; Norris, 2006; Donnarumma et al., 2017; Friston et al., 2012). Remarkably, this epistemic objective, aimed at uncertainty reduction, emerges naturally from active inference, as explained in Section 1. In summary, our model offers a comprehensive account of prediction-based written text processing, involving hierarchical inference of the content being read, prediction of upcoming textual elements, and active testing of predictions through saccades, which, in turn, inform further inference.
@@
Importantly, our model also enables the modeling of abnormal eye movements observed in reading disorders, such as dyslexia (Maisog et al., 2008; Horowitz-Kraus et al., 2014; Price, 2012). Our simulations replicate the empirical findings that dyslexics, compared to controls, exhibit fragmented text processing, characterized by an increased number of shorter saccades, both while reading single words (Simulation 1) and sentences (Simulation 2) (Zoccolotti et al., 2005; De Luca et al., 1999; Hutzler and Wimmer, 2004; Rayner, 1983). Interestingly, our model attributes these reading deficits to a straightforward mechanism: an attenuation of the role of prior information during inference (Jaffe-Dax et al., 2015). By introducing noise in the model’s transition functions, we impair its working memory of previous inference phases, forcing it to repeatedly gather evidence. Considering the cumulative effect of noise on the inference process, this mechanism accounts for why perceptual deficits in dyslexics become more pronounced for longer words and sentences (Zoccolotti et al., 2005; De Luca et al., 1999). To summarize, the model provides a mechanistic explanation of reading disorders in dyslexia, aligning well with prediction-based theories of language processing and reading.
@@
However, it is crucial to note that dyslexia is highly heterogeneous, and its underlying causes are still heavily debated (Perry et al., 2019; Zorzi et al., 2012; Ramus, 2003; Stein and Walsh, 1997; Tallal, 2004). While the attenuation of priors (Jaffe-Dax et al., 2015) explored in this article may not account for all aspects of dyslexia, our model is easily extendable to incorporate other (non-alternative) mechanisms that could contribute to reading impairments. For instance, various proposals point to disorders of low-level information sampling and attention shifting (Hari et al., 2001), or a visual attention span disorder characterized by a reduction in the number of letters that can be processed in parallel (Prado et al., 2007). These proposals (and others) can be readily integrated into the model by modifying the likelihood function that maps letters to syllables, as described in Section 1. Future research could utilize different variants of the presented model (e.g., with different levels of noise in the likelihood function and/or the transition function) to compare alternative proposals side-by-side, shedding light on the mechanisms of dyslexia and other reading difficulties, and informing the development of effective interventions to improve reading skills.

### v3

@@ 3.2 Simulation 2: Reading Sentences
This task involves reading and recognizing 100 sentences composed of 4 or 8 words (all of which are included in the dictionary of BERT (Devlin20194171, )), with each word consisting of 1 to 4 syllables. The sentences with 4 words have an average length of $25.96\pm 1.96$ letters within the range of $[21,31]$. Similarly, sentences with 8 words have an average length of $50.24\pm 1.64$ letters, in the interval of $[47,55]$ (see Table S.13). To make the text challenging, we carefully designed sentences with a substantial word overlap (Table S.15). For this task, we utilize the comprehensive model depicted in Figure 1, encompassing three hierarchical levels. At the beginning of the simulation, we use BERT to generate prior probability distributions over upcoming sentences, comprising the next 4 or 8 words. These distributions are continuously updated during reading as new syllables and words are inferred through informative saccades.
Previous studies have shown that proficient readers can scan lines of text using only a few saccades, while dyslexic individuals exhibit an increased number of shorter saccades (hutzler2004eye, ; rayner1983eye, ; de1999eye, ) and fewer word skipping occurrences (bucci2008poor, ; jainta2011dyslexic, ; hawelka2010dual, ) compared to controls. Additionally, their performance decrease and increased number of saccades are influenced by the number of words in the text (see (de1999eye, ), Fig. 2). To verify whether our model accurately replicates these findings, we compare a Control model (CM) with no noise and a Dyslexic model (DM) with noise introduced in the transition functions of all three hierarchical levels (setting the parameter $\delta^{(1)}=\delta^{(2)}=\delta^{(3)}=0.15$, see Methods section).
Figure 7: Simulation 2 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at both hierarchical levels, and three alternative versions of the Dyslexic Model, where we added noise only at the level of syllables (DM - Noise on Level 1), words (DM - Noise on Level 2), or sentences (DM - Noise on Level 3). See Table S.7 and Table S.8 for detailed results and statistical comparisons.
@@
Collectively, these results build upon those observed for single words (Simulation 1) and successfully replicate empirical findings on dyslexic individuals’ reading of sentences. The trends illustrated in Figure 7(a) closely resemble the outcomes reported (de1999eye, , Fig. 3(A)), which demonstrates that dyslexics make more saccades compared to controls while reading single lines of text (equivalent to reading 8-word sentences in Figure 7(a)). Moreover, the number of saccades exhibited by the DM in our simulations aligns well with the numerical outcomes from the same study. Similarly, the outcomes in Figure 8(a) closely resemble the findings depicted in (de1999eye, , Fig. 3(B)), revealing that dyslexics make shorter forward saccades than controls. In this context, the amplitudes of both dyslexics and the DM model closely correspond. However, the larger amplitude of backward saccades observed in our DM (Figure 8b) is in contrast with the shorter amplitude of regressions observed in dyslexic readers in (de1999eye, , Fig. 3d).
Figure 8: Results of Simulation 2: amplitude of forward (a) and backward (b) saccades when reading 4-word and 8-word sentences. The figure compares the Control Model (CM) with no noise and the Dyslexic Model (DM) with noise at all the hierarchical levels. See Table S.9 and Table S.10 for detailed results and statistical comparisons.
In summary, our results successfully replicate the observation that dyslexic individuals tend to produce a higher number of shorter saccades while reading sentences, and that their reading performance is influenced by the number of words in the sentence (see (de1999eye, , Fig. 2)). The underlying causes of these impairments are similar to those identified in the single-word reading task (Simulation 1), but their effects are amplified in sentence reading, where comprehension depends on integrating information over longer timescales. This also helps explain why noise at the sentence level significantly disrupts task performance (Figure 7). While these qualitative results are promising, further work is needed to achieve a closer—and more quantitative—alignment between the model and empirical findings on dyslexia, as well as to calibrate model parameters appropriately. For instance, our DM generates a relatively high number of regressive (backward) saccades. This occurs because the model often loses contextual information and must backtrack to earlier words in the sentence to reestablish it. However, this pattern is not commonly observed in dyslexic readers ((de1999eye, , Fig. 3d)), and the reasons for this discrepancy remain to be explored.
@@
Up to this point, our simulations involved reading known words and sentences, already present in the BERT vocabulary. Hence, the model could confidently assign probabilities at the second and third layers. Nevertheless, our model can also read unknown words and sentences, despite its inability to assign them a prior probability. When confronted with unknown words or sentences, the model reads syllable-by-syllable (or word-by-word) and simultaneously acknowledges uncertainty at the second (or third) level. This mode aligns with the nonlexical (or sublexical) route in dual-route theories of reading (coltheart2005modeling, ).
We conducted a simulation consisting of 100 trials, each involving the reading and recognition of a four-word sentence, designed for comparison with the corresponding tests in Simulation 2. In each trial, one of the four words was unknown—specifically, it was removed from the model’s vocabulary. This resulted in a set of 100 unique unknown words, each ranging from 1 to 4 syllables in length (see Table S.16 for the complete list). The sentences had an average length of $25.46\pm 1.92$ letters, with a range of $[22,32]$. A two-sample t-test comparing sentence lengths between this set and the set with only known words used in Simulation 2 revealed no significant difference ($p\approx 0.27$), confirming that the two sets were well matched in terms of both word and sentence length. Trials were terminated upon either successful completion of the sentence or failure of the model to recognize a word. The simulation results indicate that the total number of saccades increased substantially when reading sentences containing unknown words (mean $\approx 12.04$) compared to Simulation 2 (mean $\approx 4.36$); see Table 1 for statistical comparisons. Furthermore, while sentence recognition at Level 3 in Simulation 2 was nearly perfect ($p(C^{3}|O)\approx 1$), it dropped to approximately $0.49$ (on average) in the presence of unknown words, indicating a failure to correctly recognize the sentence.
Table 1: Simulation 3. Comparison of two conditions: reading 100 sentences from Simulation 2 (4 words – all known) versus 100 sentences in which one word was unknown, i.e., removed from the model’s vocabulary (4 words – 1 unknown). The table summarizes the comparison between these two conditions, reporting the mean, standard error, and p-value of t-tests for three key measures: sentence length, total number of saccades during reading, and the probability of sentence recognition.
Sentence Length Total number of Saccades Probability
4 words – Known $25.96\pm 1.96$ $4.36\pm 0.11$ $1.00$ (SE $10^{-15}$)
4 words – 1 Unknown $25.46\pm 1.92$ $12.04\pm 0.44$ $0.49\pm 0.04$
p-value $p\approx 0.27$ $p<0.001$ $p<0.001$
@@
To illustrate the model’s behavior, we describe a representative example involving the reading of the final four words of the sentence This paper is also framed in an offbeat manner, comparing two versions of the model: one in which all the words exist in the model’s vocabulary (Figure 9(a)) and another in which the word offbeat was removed from the vocabulary (Figure 9(b)). In the version with all known words, the model behaves similarly to the previous simulations, accurately recognizing all the words and the sentence. This is evident in Figure 9(a), where probabilities at Levels 2 and 3 quickly converge to one after a few iterations.
@@
Conversely, the version without knowledge of the word offbeat repeatedly makes saccades to the syllables of this word (at Level 1) until it reaches the maximum number of iterations set for this simulation. It correctly recognizes each syllable and could therefore potentially read the word aloud “non-lexically” (and potentially add it to its vocabulary) (coltheart2005modeling, ). However, it does not recognize the word (at Level 2) or the sentence (at Level 3). As shown in Figure 9(b), both the probability of the unknown word at Level 2 and the probability of the sentence at Level 3 remain below 0.5, indicating a failure of accurate recognition. After failing to recognize the unknown word (offbeat), the Level 2 inference settles on the most probable alternative in the model’s vocabulary (e.g., upbeat). Unlike the simulation described earlier, here the inference process continues even after an unrecognized word. As a result, the model is able to correctly identify the final word in the sentence (manner).
@@
The example in Fig. 9 illustrates the distinction between reading a known versus a novel word and the termination conditions for inference at each hierarchical level. The inference can terminate in two ways: first, when the model reaches a threshold of confidence about the to-be-recognized syllable, word, or sentence (i.e., the Expected Ambiguity over hidden states of Equation 5 falls below a fixed threshold, $\chi^{(i)}=1/8$). This happens when the model successfully infers a known syllable, word, or sentence.
@@
As shown in Figure 9, Level 1 successfully infers all the syllables within 3 iterations and never reaches $K_{\max}^{(1)}$. However, Level 2 fails to recognize the unknown word offbeat: it continues jumping between the syllables off and beat until it reaches the maximum number of allowed iterations (6 iterations, indicated in the grey panel). When the Level 2 inference of the current word is halted, a low confidence message is reported to the level above while inference may still proceed with the next words (that are successfully recognized). Similarly, Level 3 fails to infer the unknown sentence after the maximum number of iterations (7 iterations), despite correctly recognizing most of the words. At the end of the simulation, the model identifies that the cause of failure is at Level 2, as Level 1 correctly recognized all the syllables. Consequently, the model reads the unknown word syllable-by-syllable, i.e. using the nonlexical route in dual route theories of reading (coltheart2005modeling, ).
Figure 9: Simulation 3 results: Reading known vs. unknown words. Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time, while reading the final four words of the sentence This paper is also framed in an offbeat manner, in two cases: (a) when all the words are known and (b) when the word offbeat is unknown (i.e., removed from the vocabulary). The blue circles at the bottom level indicate saccades. The figure format is consistent with Figure 1(d).
@@ 3.4 Simulation 4: Knowing the topic improves sentence recognition
To illustrate this, we conduct a task involving the recognition of 100 sentences, each containing 9 words, generated by BERT. These sentences comprise an average of $69.55\pm 5.02$ letters, ranging between 61 and 82 characters (see Table S.13). We allocate one-third of the sentences to each of the three topics of the European Research Council (ERC): Physical Sciences and Engineering (PE), Life Sciences (LS), and Social Sciences and Humanities (SH); see Table S.17 for a full specification of the sentences and their respective topics. Please note that in this simulation, we only use topics at the third level.
We compare the performance of two versions of the model: one with a flat prior (33.3%) regarding the sentence being read, and another with an informative prior that assigns a probability of 100% to the correct topic. As depicted in Figure 10, the model with an informative prior requires significantly fewer total saccades and backward saccades than the model with a flat prior (see Table S.11 and Table S.12 for statistical comparisons). This advantage arises because knowing the topic allows the model to predict certain words with high confidence, enabling it to skip unnecessary saccades during the reading process. Intriguingly, despite the model with an informative prior skipping several words, it does not necessitate more backward (corrective) saccades compared to the model with a flat prior. This indicates that the prior information promotes a better speed-accuracy trade-off.
Figure 10: Simulation 4 results: How prior information affects the reading process. The figure displays the total saccades (a) and backward saccades (b) of two model versions: one with a flat prior about the topic of the text (flat prior) and another with certain knowledge of the topic (informative prior). The comparison highlights the impact of prior information on the number and nature of saccades during reading. See Table S.11 and Table S.12 for detailed results and statistical comparisons.
@@ 4 Discussion
In this paper, we presented a novel hierarchical active inference model of reading, which combines the robust generative abilities of large language models (Devlin20194171, ; radford2019language, ) with the inferential and information-gathering capacities of active inference (friston2010free, ; parr2022active, ).
@@
Large language models, like BERT (Devlin20194171, ) and GPT (Generative Pre-trained Transformer) (radford2019language, ), have demonstrated remarkable performance in various natural language processing tasks, including question answering, language translation, text classification, and sentence prediction. These models can learn essential linguistic structures without external supervision, making them valuable tools to explore language processing in the brain (manning2020emergent, ; goldstein2022shared, ). However, they lack the active strategies employed by human readers, such as the ability to make saccades selectively on the most informative parts of the text (ferro2010reading, ; norris2006bayesian, ).
@@
To capture active reading, we integrate BERT (Devlin20194171, ) into a hierarchical active inference framework (friston2010free, ). Active inference is a theoretical framework that describes how the brain employs probabilistic inference over a generative model to effectively sample sensory inputs, minimizing variational free energy or prediction errors – and that has been applied too many tasks, such as perceptual processing, goal-directed navigation, robot control, and social interaction (hwang2018dealing, ; maisto2023interactive, ; oliver2021empirical, ; van2023bridging, ; priorelli2023modeling, ; parr2022active, ). For modeling reading, we utilize a hierarchical generative model, aligning with evidence suggesting that language processing in the brain operates through hierarchically organized predictions and prediction errors (schmitt2021predicting, ; caucheteux2023evidence, ; heilbron2022hierarchy, ).
@@
By placing BERT at the top of the hierarchy, we ensure accurate next-word predictions and the model’s capability to tackle real-world reading tasks. Beyond next-word prediction, our hierarchical model facilitates both inferring the content being read and predicting future elements across different linguistic levels, such as syllables, words, and sentences. Our simulation results demonstrate the model’s perfect accuracy in inferring words (Simulation 1) and sentences (Simulation 2) during reading. Additionally, this hierarchical structure enables the model to read novel words not present in BERT’s vocabulary (Simulation 3), assembling them syllable-by-syllable, representing the nonlexical (or sublexical) route in dual route theories of reading (coltheart2005modeling, ). Moreover, the model can incorporate prior knowledge, like the topic of the sentence, to expedite the reading process (Simulation 4). Significantly, the model offers interpretability through its alignment of distinct levels with established linguistic constituents (sentences, words, syllables), which could correspond to distributed neural populations or cell assemblies within the brain (heilbron2022hierarchy, ; pulvermuller1999words, ). While delving into the neural mechanisms that underlie this proposed framework extends beyond the confines of this article, future studies could encompass the simulation of neuronal dynamics during the inference process, in terms of variational free energy minimization (friston2017active, ; isomura2022experimental, ).
@@
Using active inference, our model characterizes reading as an active, hypothesis testing process, allowing us to simulate eye movements during reading. The core idea is that saccades are directed to the most informative parts of the text, enabling the model to test its predictions and reduce uncertainty about the text being read (ferro2010reading, ; norris2006bayesian, ; donnarumma2017action, ; friston2012perceptions, ). Remarkably, this epistemic objective, aimed at uncertainty reduction, emerges naturally from active inference, as explained in Section 2. In summary, our model offers a comprehensive account of prediction-based written text processing, involving hierarchical inference of the content being read, prediction of upcoming textual elements, and active testing of predictions through saccades, which, in turn, inform further inference.
@@
Interestingly, our model also enables the simulation of abnormal eye movement patterns observed in reading disorders such as dyslexia (maisog2008meta, ; horowitz2014reading, ; price2012review, ). While the simulations presented here serve as a proof of concept, they qualitatively replicate empirical findings showing that individuals with dyslexia, compared to neurotypical readers, exhibit fragmented text processing. This is characterized by an increased number of shorter saccades during the reading of both single words (Simulation 1) and sentences (Simulation 2) (zoccolotti2005word, ; de1999eye, ; hutzler2004eye, ; rayner1983eye, ). Our model attributes these reading deficits to a specific disorder of predictive processing: namely, an attenuation of the influence of prior information during inference (jaffe2015computational, ). By introducing noise into the model’s transition functions, we impair its working memory of prior inference steps, thereby forcing it to repeatedly gather evidence. This mechanism, which reflects the cumulative impact of noise on inference, accounts for why perceptual difficulties in dyslexia become increasingly pronounced with longer words and more complex sentence structures (zoccolotti2005word, ; de1999eye, ).
@@
However, it is crucial to note that our simulations only aim to provide a proof of concept that disorders of reading such as dyslexia can be aligned with hierarchical predictive processing theories. Dyslexia is highly heterogeneous, and its underlying causes are still heavily debated (perry2019understanding, ; zorzi2012extra, ; ramus2003developmental, ; stein1997see, ; tallal2004improving, ; ziegler2020learning, ). It is unlikely that the attenuation of priors (jaffe2015computational, ) explored in this article would provide a comprehensive account of all aspects of dyslexia. However, our model is easily extendable to incorporate other (non-alternative) mechanisms that could contribute to reading impairments. For instance, various proposals point to disorders of low-level information sampling and attention shifting (hari2001left, ), or a visual attention span disorder characterized by a reduction in the number of letters that can be processed in parallel (prado2007eye, ). Additionally, errors in eye movements (overshooting and undershooting) can be incorporated, by introducing noise in the (transition) function mapping eye movements to subsequent spatial positions. These proposals (and others) can be readily integrated into the model by modifying the likelihood function that maps letters to syllables, as described in Section 2. Furthermore, while this study used predefined levels of noise in the transition models, these parameters can be fit to human data. Such extensions may pave the way for more comprehensive investigations of dyslexia within the predictive processing framework.
A limitation of this study is that for simplicity, it uses a relatively simple LLM: BERT (Devlin20194171, ). We selected BERT because it is computationally efficient and provides easy access to explicit probability distributions over the next words and sentences. Future work could adopt more advanced LLM models, to provide more robust linguistic predictions. Furthermore, we made some simplifying assumptions. For example, we assumed that syllables, words and sentences are represented as sequence of letters and that reading uses an absolute letter position coding. Future work could investigate alternative assumptions, see for example (snell2022relative, ) for a discussion of the absolute versus relative position coding in reading. Furthermore, future work could replace the the priors over policies that make transitions between word locations illustrated in Figure 2 with a more realistic distribution derived from empirical data. Future studies could also explore more systematically the role of the different model parameters and hyperparameters (e.g., word and sentence length, amount of noise) and address more complex linguistic datasets.
Additionally, future work could compare side by side the model presented here with established models in the field, such as connectionist personalized models, which have been very successful in explaining deficits observed in dyslexia (perry2019understanding, ; ziegler2020learning, ), and models that simulate eye movements, such as E-Z Reader (reichle1998toward, ), SWIFT (engbert2005swift, ; seelig2020bayesian, ; rabe2021bayesian, ), Über-Reader (reichle2021computational, ; veldre2020towards, ), Glenmore (reilly2002glenmore, ; reilly2006some, ), SEAM (rabe2024seam, ), the rational model of eye movements (bicknell2010rational, ), and OB1-Reader (snell2018ob1, ). By fitting specific parameters of the presented model (e.g., different levels of noise in the transition and/or likelihood functions), it would be possible to develop personalized models (or computational phenotypes (schwartenbeck2016computational, )), potentially shedding light on the mechanisms underlying reading difficulties and informing the development of effective interventions to improve reading skills.
The model presented here could be expanded to capture hierarchically higher and more abstract aspects of language processing, such as those related to shared narratives, frames and scripts, which characterize our cultural niches (minsky1986society, ; schank2013scripts, ). When opportunely placed within hierarchical architectures, these higher-level constructs could produce a cascade of predictions that contextualize and guide lower-level linguistic predictions and eye movements (bouizegarene2024narrative, ; friston2024designing, ; pezzulo2025predictive, ). The presence of these additional constructs creates a complex interplay between prior predictions at different hierarchical levels – for example, those induced by scripts, statistical regularities of sentences and word frequencies – and between prediction errors arising at different levels, which remain to be investigated in future studies.
In principle, the model presented here could also be extended to investigate developmental aspects of language and reading acquisition. Using a pretrained LLM such as BERT can be interpreted as mimicking the capabilities of adult readers. To simulate developmental trajectories, one could impose constraints on BERT’s vocabulary or its ability to generate accurate sentence completions. Limiting the vocabulary would force the model to rely more heavily on syllable-by-syllable reading, engaging a non-lexical processing route. Reducing prediction accuracy would result in less confident priors, leading the model to skip fewer words, commit more reading errors, and engage in more fixations and backward saccades. Whether a model with these constraints can successfully replicate the patterns observed during reading development remains to be evaluated in future studies.
Another important direction for future research is the integration of more realistic perceptual modules. This would allow the model to capture the complexity of the feature extraction process (e.g., (testolin2017letter, )), the parallel processing of multiple letters (e.g., (hannagan2021emergence, )), and other perceptual factors known to influence dyslexia, such as the beneficial effects of increased letter spacing (e.g., (zorzi2012extra, )). A possible approach to realize this integration is to employ neural networks as front-end perceptual modules that provide probabilistic outputs (e.g., categorical distributions over letters or syllables), which can then be incorporated as likelihoods within a Bayesian model (george2017generative, ; sancaktar2020end, ; priorelli2023flexible, ).
Finally, it is worth noting that the hierarchical control architecture proposed here can be extended to a wide range of cognitive tasks beyond reading. Many such tasks can be naturally decomposed in hierarchical control terms by considering the separation of timescales between slower, higher-level perceptual and action processes and faster, lower-level ones (botvinick2009hierarchically, ; pezzulo2018hierarchical, ; koechlin2007information, ; friston2008hierarchical, ). For instance, spatial navigation often involves high-level planning over goals or subgoals and low-level motor actions required to reach them (van2023bridging, ; donnarumma2016problem, ; tomov2020discovery, ; donnarumma2017action, ). Similarly, action observation and understanding can be structured hierarchically by distinguishing between distal goals, proximal intentions, and immediate motor acts (proietti2023active, ). Future work could adapt the hierarchical framework introduced here to explore these and other cognitive domains, shedding light on how flexible, multi-timescale processing supports complex behavior.